OpenStudy (anonymous):

the ordered pair is (-1,1) is that a solution of -4x+2y=2 x+y=-2 yes or no?

OpenStudy (anonymous):

okay so in order to find that substitute -1 into x and if y comes out to be 1 then yes it's true but it should be done for both equations inorder to find out

OpenStudy (anonymous):

I do not think so and here is why. When you manipulate the first equation you should get an equation that looks like y = (2+4x)/2 which is the same as y = 2x +1 in simplified form. The second equation is y=-2-x. If you plug in -1 into the x's of both of these equations you get y = -1 NOT 1. So no I do not think the ordered pair for these equations is (-1, 1) rather the correct ordered pair is (-1, -1).

OpenStudy (mathmale):

You have at least two choices here: 1) You could substitute x=-1 and y=1 into both equations separately; if both equations are then true, then (-1,1) is a solution of the system of linear equations -4x+2y=2 x+y=-2 OR You could solve the system of linear equations -4x+2y=2 x+y=-2 and determine whether or not the given (-1,1) matches your solution. Which will it be?

OpenStudy (anonymous):

so nope